# LLM-Based Emulation of the Radio Resource Control Layer: Towards AI-Native RAN Protocols

### v3

@@ I Introduction
Large AI Models (LAMs) are poised to transform wireless networks by enabling autonomous, cognitive capabilities that go beyond the scope of traditional optimization methods[gao2025enabling]. The forthcoming sixth generation (6G) of mobile networks is expected to incorporate Artificial Intelligence (AI) methods not merely as auxiliary optimisation tools but also as integral design primitives of the AI‑Native Air Interface (AI-AI). Recent proposals for an AI-AI posit that Machine Learning (ML) models could co‑design Physical Layer (PHY) and Medium Access Control (MAC) procedures [pl_1], thereby enabling radio stacks that adapt automatically to dynamic service requirements, resource constraints, and propagation conditions [Hoydis2021AI]. The rapid evolution of LAMs [OpenAI2025GPT4o, Google2023Gemini] strengthens this vision—when endowed with multi-modal perception and reasoning, such models become plausible agents for protocol‑centric tasks that demand both linguistic competence and structured decision making. This raises a timely research challenge: determining the extent to which LAMs can directly engage in the generation and interpretation of control signaling in the Radio Access Network (RAN). Can LAMs achieve sufficient protocol literacy for RRC [3gpp38331] emulation? What fine-tuning strategies are effective? What is the performance fidelity of such a system? These are the research questions this paper intends to answer. We hypothesize that LAM‑scale pretraining provides the compositional priors needed to generalize across 3GPP releases and vendor variants, and to perform multi‑procedure, stateful reasoning under incomplete context—capabilities that compact task‑specific transformers or grammar‑only decoders lack without extensive bespoke engineering.
@@
Without loss of generality, this work focuses on decoder‑only transformer architectures (e.g., Llama‑class models [Touvron2023LLaMA]). Autoregressive decoding matches the turn‑taking structure of RRC exchanges and permits key–value caching, which curtails inference latency and memory footprint, both critical for deployment at the gNB. Decoder‑only models thereby strike a favourable balance between capacity and computational efficiency, especially when combined with parameter‑efficient fine‑tuning techniques such as LoRA [Hu2021LoRA]. LoRA freezes the original model weights and introduces low‑rank update matrices, reducing trainable parameters by orders of magnitude while maintaining expressiveness; the approach mitigates both the cost of domain adaptation and the runtime overhead of inference, two primary barriers to operational use.
Training LAMs for specialist telecom domains demands curated datasets that capture the combinatorial breadth of protocol states, while deployment must meet stringent latency budgets to avoid impairing control‑plane responsiveness. To address these issues, we rebuilt two data corpuses of RRC traces (4G and 5G) with multi-operator coverage. Then, we adopt a standards-aware evaluation suite that combines ASN.1 [ITUT_ASN1_Intro] syntax validation, schema-level field coverage, and UL→DL state-machine conformance, reporting these jointly with semantic similarity and latency across 120 configurations. Ultimately, we also outline a forward-looking engineering plan toward sub-100 ms inference latency via quantization [QLORA], constrained decoding, KV caching and multi‑token draft‑verify decoding.
@@ III-A Datasets: Construction, Scope, and Coverage
Both corpora are private and proprietary. Public LTE datasets/simulators exist (e.g., [liu_rrc_2025, liu_nrRRC_simulator_2025]), but are synthetic and were not used in this work.
@@ III-C Training Configuration and Convergence Behavior
Learning Rate $2\times 10^{-5}$
$2\times 10^{-5}$ (full FT)
$3\times 10^{-4}$ (LoRA)
@@
LR schedule
Cosine decay
+ 100-step warmup
@@ V Conclusion and Future Work
    Inference acceleration: speculative multi-token decoding [leviathan2022specdec], hierarchical agent architectures, and lightweight backbones to reconcile fidelity with real-time constraints.
@@
    Extended-context modeling: the context window length of current LLM s is critical for cross-layer reasoning and complex RRC procedures. Longer contexts or memory-augmented transformers (e.g., RoPE-extended variants, ALiBi [press2022alibi]) will be explored to eliminate segmentation and preserve procedural continuity.

### v4

@@ I Introduction
Large AI Models (LAMs) are poised to transform wireless networks by enabling autonomous, cognitive capabilities that go beyond the scope of traditional optimization methods[3]. The forthcoming sixth generation (6G) of mobile networks is expected to incorporate Artificial Intelligence (AI) methods not merely as auxiliary optimisation tools but also as integral design primitives of the AI‑Native Air Interface (AI-AI). Recent proposals for an AI-AI posit that Machine Learning (ML) models could co‑design Physical Layer (PHY) and Medium Access Control (MAC) procedures [15], thereby enabling radio stacks that adapt automatically to dynamic service requirements, resource constraints, and propagation conditions [4]. The rapid evolution of LAMs [10, 13] strengthens this vision—when endowed with multi-modal perception and reasoning, such models become plausible agents for protocol‑centric tasks that demand both linguistic competence and structured decision making. This raises a timely research challenge: determining the extent to which LAMs can directly engage in the generation and interpretation of control signaling in the Radio Access Network (RAN). Can LAMs achieve sufficient protocol literacy for RRC [1] emulation? What fine-tuning strategies are effective? What is the performance fidelity of such a system? These are the research questions this paper intends to answer. We hypothesize that LAM‑scale pretraining provides the compositional priors needed to generalize across 3GPP releases and vendor variants, and to perform multi‑procedure, stateful reasoning under incomplete context—capabilities that compact task‑specific transformers or grammar‑only decoders lack without extensive bespoke engineering.
@@
Without loss of generality, this work focuses on decoder‑only transformer architectures (e.g., Llama‑class models [14]). Autoregressive decoding matches the turn‑taking structure of RRC exchanges and permits key–value caching, which curtails inference latency and memory footprint, both critical for deployment at the gNB. Decoder‑only models thereby strike a favourable balance between capacity and computational efficiency, especially when combined with parameter‑efficient fine‑tuning techniques such as LoRA [5]. LoRA freezes the original model weights and introduces low‑rank update matrices, reducing trainable parameters by orders of magnitude while maintaining expressiveness; the approach mitigates both the cost of domain adaptation and the runtime overhead of inference, two primary barriers to operational use.
Training LAMs for specialist telecom domains demands curated datasets that capture the combinatorial breadth of protocol states, while deployment must meet stringent latency budgets to avoid impairing control‑plane responsiveness. To address these issues, we rebuilt two data corpuses of RRC traces (4G and 5G) with multi-operator coverage. Then, we adopt a standards-aware evaluation suite that combines ASN.1 [6] syntax validation, schema-level field coverage, and UL→DL state-machine conformance, reporting these jointly with semantic similarity and latency across 120 configurations. Ultimately, we also outline a forward-looking engineering plan toward sub-100 ms inference latency via quantization [2], constrained decoding, KV caching and multi‑token draft‑verify decoding.
@@ III-A Datasets: Construction, Scope, and Coverage
Both corpora are private and proprietary. Public LTE datasets/simulators exist (e.g., [9, 8]), but are synthetic and were not used in this work.
@@ III-C Training Configuration and Convergence Behavior
Learning Rate $2\times 10^{-5}$ $2\times 10^{-5}$ (full FT) $3\times 10^{-4}$ (LoRA)
@@
LR schedule Cosine decay + 100-step warmup
@@ V Conclusion and Future Work
    Inference acceleration: speculative multi-token decoding [7], hierarchical agent architectures, and lightweight backbones to reconcile fidelity with real-time constraints.
@@
    Extended-context modeling: the context window length of current LLM s is critical for cross-layer reasoning and complex RRC procedures. Longer contexts or memory-augmented transformers (e.g., RoPE-extended variants, ALiBi [11]) will be explored to eliminate segmentation and preserve procedural continuity.